# The Efficacy of Semantics-Preserving Transformations in Self-Supervised Learning for Medical Ultrasound

### v1

@@ Appendix G Additional Positive Pair Examples
Figs. 1011, and 12 provide several examples of positive pairs produced by the BYOL, AugUS-O, and AugUS-D pipelines, respectively. Each figure shows original images from LUSData, along with two views of each image that were produced by applying stochastic data augmentation twice to the original images.
Figure 10: Examples of lung ultrasound images (left) and positive pairs produced using the BYOL pipeline (right).
@@
Appendix H Label Efficiency Statistical Testing
@@
For each of the AB and PE tasks, there were five experimental conditions: SimCLR pretraining with the BYOL pipeline, SimCLR pretraining with the AugUS-O pipeline, SimCLR pretraining with the AugUS-D pipeline, ImageNet weight initializations, and random weight initialization. The population consisted of $20$ subsets of the training set that were split randomly by patient. The same splits were used across all conditions, reflecting a repeated measures design. The experiment was repeated separately for the AB and the PE task.
@@
The Friedman Test Statistic ($F_{r}$) was $75.44$ for AB, with a p-value of 00. For PE, the $F_{r}=45.44$, and the p-value was 00. As such, the null hypothesis was rejected for both cases, indicating the existence of differences among the mean test AUC across conditions. The Wilcoxon Sign-Rank test was performed as a post-hoc test between each pair of populations. Tables XVII and XVIII provide all Wilcoxon Test Statistics, along with $p$-values and differences of the medians between conditions. The Bonferroni correction was applied to the $p$-values to keep the family-wise error rate to $\alpha=0.05$. Statistically significant comparisons are indicated.
@@
TABLE XVII: Test statistics ($T$) and p-values obtained from the Wilcoxon Sign-Rank post-hoc tests comparing LUSData test AUC on AB for classifiers trained on subsets of the training set. For comparison $a/b$, $\Delta:=\text{median}(b)-\text{median}(a)$. The displayed $p-$values have been adjusted according to the Bonferroni correction to control the family-wise error rate.
@@
TABLE XVIII: Test statistics ($T$) and p-values obtained from the Wilcoxon Sign-Rank post-hoc tests comparing LUSData test AUC on PE for classifiers trained on subsets of the training set. For comparison $a/b$, $\Delta:=\text{median}(b)-\text{median}(a)$. The displayed $p-$values have been adjusted according to the Bonferroni correction to control the family-wise error rate.
@@
Appendix I Additional Random Crop and Resize Experiments
The C&R transform encourages pretrained representations to be invariant to scale. It is also believed that the C&R transform instills invariance between global and local views or between disjoint views of the same object type [30]. While the minimum area of the crop determines the magnitude of the scaling transformations, the aspect ratio range dictates the difference in distortion in both axes of the image. The default aspect ratio range is $[0.75,1.\overline{33}]$. We pretrained with the AugUS-D pipeline using a fixed aspect range of $1$ and $c=0.08$, which resulted in test AUC of $0.971$ for AB and $0.881$ for PE. Compared to the regular AugUS-D that uses the default aspect ratio range (Table 5), AB test AUC is unchanged; however, PE test AUC decreased by $0.016$.
Lastly, we conducted pretraining on LUSData using only the C&R transformation; that is, the data augmentation pipeline was [B00]. Recent work by Moutakanni et al. [41] suggests that, with sufficient quantities of training data, competitive performance in downstream computer vision tasks can be achieved using crop and resize as the sole transformation in joint embedding SSL. Linear evaluation of a feature extractor pretrained with only C&R yielded test AUC of $0.964$ and $0.874$ on AB and PE, respectively. Compared to the linear evaluations presented in Section 4, these metrics are greater than those achieved using AugUS-O, but less than thse achieved with BYOL or AugUS-D. It is evident that C&R is a powerful transformation for detecting local objects.

### v2

@@ Appendix G Additional Positive Pair Examples
Figs. 1011, and 12 provide several examples of positive pairs produced by the StandardAug, AugUS-O, and AugUS-D pipelines, respectively. Each figure shows original images from LUSData, along with two views of each image that were produced by applying stochastic data augmentation twice to the original images.
Figure 10: Examples of lung ultrasound images (left) and positive pairs produced using the StandardAug pipeline (right).
@@
Appendix H Results with ResNet18 Backbone
As outlined in Section 3.5, MobileNetV3Small [32] was selected as the feature extractor for all experiments in this study, primarily due to its suitability for lightweight inference in edge deployment scenarios. It has also been used in prior work for similar tasks [33, 24]. However, we wanted to determine if similar trends in the results held for an alternate backbone architecture with greater capacity. Larger CNN backbones are frequently used for self-supervised pretraining.
We repeated the experiments in Sections 4.2 and 4.3 using ResNet18 [40] as the feature extractor, which is a more common architecture. Note that ResNet18’s capacity is far greater than that of MobileNetV3Small — the former has approximately 11 200 0001120000011\,200\,00011 200 000 trainable parameters, while the latter has about 927 000927000927\,000927 000. We applied the same architecture for the projector head that was used for MobileNetV3Small experiments, which was a $2$-layer multilayer perceptron with $576$ nodes comprising each layer.
Table 15 displays the performance of pretrained ResNet18 backbones when evaluated on the LUSData test set for the AB, PE, and COVID tasks. Most fine-tuned models exhibited overfitting, likely due to the architecture’s capacity. Notably, the pretrained models tended to exhibit overfitting when fine-tuned, while the fully supervised ImageNet-pretrained model achieved strong performance on AB and PE. Linear classifiers trained on frozen pretrained backbones’ output features generally achieved the strongest performance and avoided overfitting. Consistent with MobileNetV3Small, the use of cropping-based augmentation pipelines translated to greater performance on AB and PE in the linear evaluation setting. In the fine-tuning setting, the backbone pretrained using AugUS-O exhibited markedly poorer performance on AB but achieved the greatest test AUC on PE. On the COVID test set, the backbone pretrained using the AugUS-O pipeline led to the greatest average AUC in both the linear and fine-tuning settings.
The fine-tuned backbones and linear classifiers were also evaluated on the external test set for the AB and PE tasks. Again, linear classifiers with pretrained backbones mostly achieved greater performance than fine-tuned classifiers. Backbones pretrained with the StandardAug pipeline achieved the greatest external test set AUC for the AB task. However, those pretrained with AugUS-O performed comparably with StandardAug for the PE task under the linear evaluation setting but achieved the greatest AUC in the fine-tuning setting.
Two key observations were drawn from these experiments. First, the low-capacity MobileNetV3Small backbones achieved similar performance to the high-capacity ResNet18 backbones for these LUS tasks when their weights were frozen (i.e., linear classifiers). Second, with a higher-capacity backbone, linear classifiers trained on the features outputted by SSL-pretrained backbones often achieved greater performance than fine-tuning using despite the weight initialization strategy. The opposite trend was observed for backbones initialized with ImageNet-pretrained weights.
Table 15: Test set performance for linear classification (LC) and fine-tuning (FT) experiments with the COVID task using pretrained ResNet18 backbones. For COVID, metrics are averages across all four classes. The best observed metrics in each experimental setting are in boldface.
Train Setting Task Weights Pipeline Accuracy Precision Recall AUC
LC AB SimCLR StandardAug $\mathbf{0.932}$ $0.929$ $\mathbf{0.845}$ $\mathbf{0.968}$
SimCLR AugUS-O $0.916$ $0.923$ $0.792$ $0.954$
SimCLR AugUS-D $0.931$ $\mathbf{0.951}$ $0.819$ $0.966$
ImageNet - $0.878$ $0.838$ $0.749$ $0.933$
PE SimCLR StandardAug $0.781$ $0.772$ $0.777$ $\mathbf{0.879}$
SimCLR AugUS-O $\mathbf{0.794}$ $\mathbf{0.806}$ $0.759$ $0.878$
SimCLR AugUS-D $0.779$ $0.739$ $\mathbf{0.840}$ $0.873$
ImageNet - $0.732$ $0.693$ $0.799$ $0.823$
COVID SimCLR StandardAug $0.537$ $0.538$ $0.496$ $\mathbf{0.827}$
SimCLR AugUS-O $\mathbf{0.567}$ $\mathbf{0.438}$ $\mathbf{0.515}$ $0.782$
SimCLR AugUS-D $0.487$ $0.353$ $0.447$ $0.789$
ImageNet - $0.527$ $0.334$ $0.476$ $0.699$
FT AB SimCLR StandardAug $0.882$ $0.796$ $0.824$ $0.929$
SimCLR AugUS-O $0.375$ $0.328$ $\mathbf{0.990}$ $0.867$
SimCLR AugUS-D $0.498$ $0.377$ $0.979$ $0.907$
Random - $\mathbf{0.919}$ $\mathbf{0.898}$ $0.830$ $0.960$
ImageNet - $0.892$ $0.778$ $0.906$ $\mathbf{0.961}$
PE SimCLR StandardAug $0.692$ $0.643$ $0.817$ $0.795$
SimCLR AugUS-O $0.773$ $\mathbf{0.928}$ $0.576$ $\mathbf{0.887}$
SimCLR AugUS-D $0.654$ $0.601$ $\mathbf{0.852}$ $0.755$
Random - $0.720$ $0.684$ $0.784$ $0.815$
ImageNet - $\mathbf{0.778}$ $0.797$ $0.727$ $0.852$
COVID SimCLR StandardAug $0.418$ $\mathbf{0.451}$ $0.447$ $0.719$
SimCLR AugUS-O $\mathbf{0.564}$ $0.434$ $\mathbf{0.511}$ $\mathbf{0.723}$
SimCLR AugUS-D $0.431$ $0.346$ $0.376$ $0.631$
Random - $0.372$ $0.277$ $0.354$ $0.629$
ImageNet - $0.549$ $0.433$ $0.496$ $0.719$
Table 16: External test set performance for linear classifiers (LC) and fine-tuned models (FT). The best observed metrics in each experimental setting are in boldface.
Train Setting Task Initial Weights Pipeline Accuracy Precision Recall AUC
LC AB SimCLR StandardAug $\mathbf{0.762}$ $\mathbf{0.956}$ $\mathbf{0.581}$ $\mathbf{0.877}$
SimCLR AugUS-O $0.713$ $0.939$ $0.494$ $0.826$
SimCLR AugUS-D $0.739$ $0.945$ $0.542$ $0.862$
ImageNet - $0.659$ $0.904$ $0.402$ $0.785$
PE SimCLR StandardAug $\mathbf{0.797}$ $0.884$ $0.784$ $\mathbf{0.887}$
SimCLR AugUS-O $0.786$ $\mathbf{0.911}$ $0.737$ $\mathbf{0.887}$
SimCLR AugUS-D $0.772$ $0.818$ $\mathbf{0.826}$ $0.851$
ImageNet - $0.701$ $0.790$ $0.723$ $0.776$
FT AB SimCLR StandardAug $0.767$ $\mathbf{0.951}$ $0.594$ $\mathbf{0.880}$
SimCLR AugUS-O $0.597$ $0.576$ $\mathbf{0.924}$ $0.728$
SimCLR AugUS-D $0.697$ $0.666$ $0.867$ $0.816$
Random - $0.744$ $0.919$ $0.576$ $0.851$
ImageNet - $\mathbf{0.771}$ $0.855$ $0.688$ $0.857$
PE SimCLR StandardAug $0.612$ $0.791$ $0.532$ $0.700$
SimCLR AugUS-O $0.678$ $\mathbf{0.972}$ $0.509$ $\mathbf{0.897}$
SimCLR AugUS-D $0.703$ $0.752$ $\mathbf{0.797}$ $0.755$
Random - $0.743$ $0.845$ $0.731$ $0.809$
ImageNet - $\mathbf{0.764}$ $0.897$ $0.712$ $0.854$
Appendix I Label Efficiency Statistical Testing
@@
For each of the AB and PE tasks, there were five experimental conditions: SimCLR pretraining with the StandardAug pipeline, SimCLR pretraining with the AugUS-O pipeline, SimCLR pretraining with the AugUS-D pipeline, ImageNet weight initializations, and random weight initialization. The population consisted of $20$ subsets of the training set that were split randomly by patient. The same splits were used across all conditions, reflecting a repeated measures design. The experiment was repeated separately for the AB and the PE task.
@@
The Friedman Test Statistic ($F_{r}$) was $75.44$ for AB, with a p-value of 00. For PE, the $F_{r}=45.44$, and the p-value was 00. As such, the null hypothesis was rejected for both cases, indicating the existence of differences among the mean test AUC across conditions. The Wilcoxon Sign-Rank test was performed as a post-hoc test between each pair of populations. Tables 17 and 18 provide all Wilcoxon Test Statistics, along with $p$-values and differences of the medians between conditions. The Bonferroni correction was applied to the $p$-values to keep the family-wise error rate to $\alpha=0.05$. Statistically significant comparisons are indicated.
@@
Table 17: Test statistics ($T$) and p-values obtained from the Wilcoxon Sign-Rank post-hoc tests comparing LUSData test AUC on AB for classifiers trained on subsets of the training set. For comparison $a/b$, $\Delta:=\text{median}(b)-\text{median}(a)$. The displayed $p-$values have been adjusted according to the Bonferroni correction to control the family-wise error rate.
@@
Table 18: Test statistics ($T$) and p-values obtained from the Wilcoxon Sign-Rank post-hoc tests comparing LUSData test AUC on PE for classifiers trained on subsets of the training set. For comparison $a/b$, $\Delta:=\text{median}(b)-\text{median}(a)$. The displayed $p-$values have been adjusted according to the Bonferroni correction to control the family-wise error rate.
@@
Appendix J Additional Random Crop and Resize Experiments
The C&R transform encourages pretrained representations to be invariant to scale. It is also believed that the C&R transform instills invariance between global and local views or between disjoint views of the same object type [13]. While the minimum area of the crop determines the magnitude of the scaling transformations, the aspect ratio range dictates the difference in distortion in both axes of the image. The default aspect ratio range is $[0.75,1.33]$. We pretrained with the AugUS-D pipeline using a fixed aspect range of $1$ and $c=0.08$, which resulted in test AUC of $0.971$ for AB and $0.881$ for PE in the LC training setting. Compared to the regular AugUS-D that uses the default aspect ratio range (Table 5), AB test AUC remained unchanged, and PE test AUC decreased by $0.005$.
Lastly, we conducted pretraining on LUSData using only the C&R transformation; that is, the data augmentation pipeline was [B00]. Recent work by Moutakanni et al. [44] suggests that, with sufficient quantities of training data, competitive performance in downstream computer vision tasks can be achieved using crop and resize as the sole transformation in joint embedding SSL. Linear evaluation of a feature extractor pretrained with only C&R yielded test AUC of $0.964$ and $0.874$ on AB and PE, respectively. Compared to the linear evaluations presented in Section 4, these metrics are greater than those achieved using AugUS-O, but less than those achieved with StandardAug or AugUS-D. It is evident that C&R is a powerful transformation for detecting local objects.